OpenStudy (anonymous):

Consider the eddy-current braking. A square loop, with 10 cm side is shot with the velocity 10 m/s into the uniform magnetic field with magnitude 0.1 T. The field is perpendicular to the plane of the loop, and the loop starts entering magnetic field at t=0. The resistance of the loop is 1.00 Ohm and the mass is 1.0 g. Assume the loop is moving to the right along x-axis and that x(t=0)=0. Find the velocity of the loop 0.1 seconds later.

OpenStudy (anonymous):

The equation of motion is x(t)=C1+C2*exp(-K*t) where C1=v0/K C2=-v0/K and K=B^2*s^2/(m*R) I have posted a full explanation to YouTube http://youtu.be/MqfCtZQ7Qag This is the first video of a 4 video playlist Use HD quality and a larger screen for best viewing results